OpenStudy (anonymous):

Without using a calculator, solve: (3/64)-(3/80)+(2/20)

OpenStudy (anonymous):

64=4 *16 80= 5*16 20=4*5 LCD=4*5*16=320 write each fraction with LCD by appropriate multiplication\[\frac{3}{64}*\frac{5}{5}-\frac{3}{80}*\frac{4}{4}+\frac{2}{20}*\frac{16}{16}\]\[=\frac{15-12+32}{320}=\frac{35}{320}\]\[=\frac{\cancel5*7}{\cancel5*64}=\frac{7}{64}\]